OpenStudy (anonymous):

When 0.5(20x – 50y + 36) – 0.25(–100x + 40y – 12) is simplified, what is the resulting expression?

OpenStudy (anonymous):

im sorry i don't know how to help you with this... :(

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

i can recommend you ask @iGreen @butterflydreamer or @tkhunny

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

if it helps i gave you a medal

OpenStudy (phi):

I would start with this part: 0.5(20x – 50y + 36) you can "distribute" the 0.5 (which is 1/2 ) that means multiply each thing (term) inside the parens by 1/2 to multiply you write 1/2 * in front of each term

OpenStudy (phi):

can you do that?

OpenStudy (anonymous):

0.5*(20x – 50y + 36) right here?

OpenStudy (phi):

the idea is to write 0.5 in front of *each* term

OpenStudy (anonymous):

see i knew someone would come to the rescue

OpenStudy (anonymous):

10x+25y+18

OpenStudy (anonymous):

:o nice

OpenStudy (phi):

close. but you should do it in steps 0.5(20x – 50y + 36) 0.5*20*x + 0.5* -50*y + 0.5*36

OpenStudy (igreen):

Yeah, you got the sign wrong.

OpenStudy (phi):

notice you got the "y" part with the wrong sign

OpenStudy (igreen):

\(\sf 0.5 \times -50 \ne 25\)

OpenStudy (phi):

can you try again?

OpenStudy (anonymous):

i got it wrong?

OpenStudy (phi):

you got 10x+25y+18 but if you start with 0.5(20x – 50y + 36) and distribute the 0.5, like this: 0.5*20*x + 0.5* -50*y + 0.5*36 now simplify the "number" part (and leave the letters as they are)

OpenStudy (anonymous):

i dont understand why you're multiplying the 0.5 with the letter and the number

OpenStudy (phi):

first, 20x is short for 20 * x (twenty times "x") when we multiply a number times a letter, it's easy. we just write 20*x or 20x (we don't bother showing the multiply sign. ok so far?

OpenStudy (anonymous):

oh true

OpenStudy (anonymous):

so 0.5*20x=10x

OpenStudy (phi):

next, think of the parens as a "package" with "stuff" inside (20x – 50y + 36) when we write 0.5 * (20x -50y +36) think like this: I want 0.5 of the whole package in other words, 0.5 = 1/2 and 1/2 (20x -50y +36) means I want 1/2 of the package and that means 1/2 of each thing inside. following ?

OpenStudy (anonymous):

oh snap I have to go he bell rang I'm going to come after school lol

OpenStudy (anonymous):

bye